# Expanding the Resolution Boundary of Outcome-Based Imperfect-Recall Abstraction in Games with Ordered Signals

### v1

@@ 4 Games with Ordered Signals
Figure 2 illustrates the game tree with ordered signal for Leduc Hold’em. According to rules of Leduc Hold’em rules, see Appendix D.1, all subtrees rooted at chance public nodes in the second stage are identical. For the sake of simplicity in visualization, we represent the Leduc game tree as a forest of two trees. However, it is worth noting that the game with ordered signals model defined in Definition 1 is a complete tree. We can concatenate a subtree from the first stage with five subtrees from the second stage to form the complete Leduc game tree. It is important to distinguish between two types of terminal public nodes. When $\tilde{z}$ is a terminal public node with fold, it implies $\sum_{i=1}^{N}\omega_{i}(\tilde{z})=1$, indicating that only one player remains in the game. On the other hand, when $\tilde{z}$ is a terminal public node with showdown, it means that more than one player is still participating in the game, and their payoffs need to be determined through a showdown. In Leduc Hold’em, Rhode Island Hold’em, HULHE, HUNL, and the Numeral211 Hold’em introduced in the experimental section, all terminal public nodes with fold are located in the final stage. Meanwhile, the signal set is solely dependent on the stage, implying that the terminal signal set $\tilde{\Theta}$ in these games corresponds exclusively to the signal set of the final stage. Unlike the definition in Gilpin and Sandholm (2007b), signals in Leduc Hold’em, as defined in Definition 1, refer to the complete set of cards currently dealt. For example, JQ in the first stage (the first player dealt a J, the second player dealt a Q) and JQK in the second stage (the first player dealt a J, the second player dealt a Q, and the community card is K) are signals in their respective stages. In the case of JQK in the second stage, the cards dealt to each player and the individual community card are observations of the signal from their respective perspectives. That is, if $\theta=\text{JQK}$, then
$\left\{\begin{array}[]{ll}\vartheta_{1}(\theta)=\text{J*K}\\ \vartheta_{2}(\theta)=\text{Q*K}\\ \vartheta_{pub}(\theta)=\text{**K}\end{array}\right..$ (4)
We can define other games with ordered signals where signal unveiling follows various probability distributions, not just the model of sampling without replacement, as we are familiar with.
@@
Given a signal $\theta$ and a player $i\in N$, we can define the signal infoset that player $i$ cannot distinguish related to $\theta$ as $\vartheta_{i}(\theta)=\{\theta^{\prime}\in\Theta\mid O_{i}(\theta)=O_{i}(\theta^{\prime})\land O_{pub}(\theta)=O_{pub}(\theta^{\prime})\}$. For example, in Figure 2, JJQ and JQJ belong to the same signal infoset for player 1. Additionally, the concept of a signal infoset space $\varTheta=(\varTheta_{1},\dots,\varTheta_{N},\varTheta_{c})$ is introduced, where $\varTheta_{i}=\{\vartheta_{i}(\theta)|\theta\in\Theta\}$ denotes the collection of signal infosets pertinent to player $i$. A game with ordered signals is a special type of imperfect-information game. We can transform a game with ordered signals into an imperfect-information game, a process illustrated in Appendix E.
@@
Compared to the definition of games with ordered signals introduced by Gilpin and Sandholm (2007b), Definition 1 is more generalized and concise. Specifically:
@@
    The definition of stages is more flexible. In the original definition, stages were predefined non-negative integers. In Definition 1, stages are defined based on the number of chance public nodes along a path. This allows games with ordered signals to be treated as a single tree structure, eliminating the need to split them into a forest based on chance public nodes.
@@ 5 Abstraction for Games with Ordered Signals
The model of games with ordered signals divides the description of nodes in imperfect-information games into two dimensions: the signals and the public nodes. Correspondingly, it categorizes the description of infosets in imperfect-information games into signal infosets and public nodes. The infosets in an imperfect-information game that can be traced back to a common public node will have the same action space and action history sequence. As long as the signal infosets are suitable for merging together, their corresponding infosets should also be suitable for merging into the same abstracted infoset. The game with ordered signals simplifies the process of infosets abstraction in imperfect-information games, eliminating the need for pre-filtering infosets. This also implies that infosets at the same stage will have consistent abstraction results based on the signal infosets abstraction. In other words, if $\vartheta_{i}^{1}$ and $\vartheta_{i}^{2}$ are two $r$-stage signal infosets for player $i$ and can be grouped into the same abstracted signal infoset during a abstraction process, then for any public node $\tilde{h}_{i}$ corresponding to player $i$ in the $r$-stage, the infoset related to $\tilde{h}_{i}$ and $\vartheta_{i}^{1}$ will always be classified into the same abstracted infoset as the one related to $\tilde{h}_{i}$ and $\vartheta_{i}^{2}$. This simplifies the workload of infoset abstraction, even though it sacrifices some flexibility in infosets abstraction, this constraint is reasonable. We introduce the following definition to provide a clearer description of the abstraction process in games with ordered signals:
Definition 2
@@
Intuitively, one might expect that more refined abstractions would result in superior strategies. However, the discovery of abstraction pathology refutes this notion. [^2] It’s worth noting that while Waugh’s experiments focus on imperfect-information games, Leduc Hold’em itself is a form of game with ordered signals. Moreover, the information set abstraction and action abstraction performed in their experiments correspond to the signal and action abstraction of games with ordered signals , respectively. Therefore, the concept of abstraction pathology also applies to the games with ordered signals. It is crucial to stress that the issue of abstraction pathology arises from abstracting the opponent’s decision space, essentially modeling the opponent’s behavior based on a fundamental assumption: that the opponent exclusively acts within the defined abstracted decision space. However, this assumption may not always hold true. A more reasonable perspective called one-player-abstraction-perspective is that players make decisions within the abstracted decision space as a simplification to adapt to the opponent’s decisions in the complete decision space. The study conducted by Waugh et al. (2009b) demonstrated that when the opponent’s decision space is not abstracted, increasing the precision of one’s own abstraction can lead to reduced exploitability of the abstracted Nash equilibrium solution in the original game. [^3] It’s important to observe that although Waugh’s conclusion is derived within the context of imperfect-information games, while games with ordered signals represent a particular subclass of imperfect-information games. Consequently, the same conclusion remains valid when applied to games with ordered signals.
@@
Although the task of signal abstraction in games with ordered signals were not explicitly defined in prior research, earlier studies employed various rudimentary terms to describe similar problems and methods. This was particularly evident in tasks such as hand abstraction in the domain of Texas Hold’em game. For instance, Shi and Littman (2001) referred to it as bins, Billings et al. (2003) as buckets, Johanson et al. (2013) mentioned state-space abstraction, and Gilpin and Sandholm (2007b) used the term signals to describe analogous concepts. Building upon these analogous descriptions of signal abstraction tasks, numerous approaches have been proposed to solve it. Among them, Gilpin and Sandholm (2007b) and Waugh (2013) developed lossless isomorphism (LI), Johanson (2007) introduced the expected hand strength (E[HS]) algorithm, while Gilpin et al. (2007); Gilpin and Sandholm (2008) introduced the potential-aware (PA) algorithm, which later led to the derivation of the potential-aware based on the earthmover’s distance (PAEMD) algorithm introduced by Ganzfried and Sandholm (2014). Our definition of the signal abstraction task has been significantly influenced by the work of Gilpin and Sandholm (2007b) and is also more closely aligned with the essence of the hand abstraction task in Texas Hold’em-style game.
@@ 6 Potential Outcome Isomorphism
In game theory, calculating the expected payoff at each node relies on the utility backpropagated from the terminal nodes. Consequently, mainstream signal abstraction algorithms cluster signals based on the showdown outcomes in the terminal stage, such as E[HS], PA, and PAEMD. These algorithms are referred to as outcome-based imperfect-recall algorithms. They possess the flexibility to adjust the scale of abstracted infosets. Our inquiry pertains to the resolution boundary, which refers to the quantity of signal infosets that the algorithm can recognize without exceeding, of such outcome-based imperfect-recall signal abstraction algorithms and whether they can achieve LI. The objective of this section is to develop a signal abstraction, termed as potential outcome isomorphism (POI), which aims to identify as many abstracted signal infosets as possible based on future showdown outcomes only. In the next section, we will elucidate why the POI can represent the resolution boundary of outcome-based imperfect-recall algorithms.
@@
2:$r\in\{1,\dots,|\mathfrak{r}|\}$. Stages.
@@
5:$\mathcal{D}_{i}^{(r+1)}:\mathbb{N}\mapsto\mathbb{N}$. Next stage potential outcome isomorphism map, if $r=|\mathfrak{r}|$, $\mathcal{D}_{i}^{(r+1)}=\emptyset$.
6:procedure StagePotentialOutcomeIsomorphism($r$, $\varTheta_{i}^{(r)}$, $\mathcal{D}_{i}^{(r+1)}$)
@@
The construction of POI, as shown in Algorithm 1, is described as follows. The algorithm begins at the last stage, $r=|\mathfrak{r}|$, where $\mathfrak{r}=\{r=\gamma(\tilde{x})\mid\tilde{x}\in\tilde{X}\}$, and proceeds bottom-up, computing each stage down to $r=1$. For any player $i$ and signal information set $\vartheta$ at stage $r$, we construct a potential outcome feature $f_{i}^{(r)}(\vartheta)$. These feature vectors consist of natural numbers and have varying dimensions across stages but remain consistent within the same stage.
After removing duplicates from the potential outcome features of all signal infosets for player $i$ at stage $r$, we form a set denoted as $\mathcal{C}_{i}^{(r)}$. Upon sorting the potential outcome features in $\mathcal{C}_{i}^{(r)}$ in lexicographical order, a unique lexicographical identifier in the range of 00 to $|\mathcal{C}_{i}^{(r)}|-1$ is assigned to each potential outcome feature $f_{i}^{(r)}$. This identifier is referred to as the lexicographical order identifier ($lexid$) for $f_{i}^{(r)}$. If two signal information sets, $\vartheta_{1}$ and $\vartheta_{2}$, in $\varTheta_{i}^{(r)}$ share the same potential outcome feature $f_{i}^{(r)}$, they are categorized as the same abstracted signal infoset.
For player $i$ at the last stage, $r=|\mathfrak{r}|$, the potential outcome feature for a signal infoset $\vartheta\in\varTheta_{i}^{(r)}$ is defined as
@@
$f_{i}^{(r),l}(\vartheta)=\left\{\begin{array}[]{ll}\sum\limits_{\theta\in\vartheta}\mathds{1}\left\{\exists j\neq i,\succeq(\theta,i,j)=\text{False and }\succeq(\theta,j,i)=\text{True}\right\}&\text{if }l=0,\\ \sum\limits_{\theta\in\vartheta}\mathds{1}\left\{\sum\limits_{j\neq i}\mathds{1}\left\{\succeq(\theta,j,i)=\text{True}\right\}=l\text{ and }\forall j\neq i,\succeq(\theta,i,j)=\text{True}\right\}&\text{if }l\neq 0.\end{array}\right.$ (8)
@@
  • $f_{i}^{(r),l}(\vartheta)$ represents the number of signals in $\vartheta$ where player $i$ is ranked no less than all other players and is ranked exactly $l$ times higher than the other players.
@@
Specifically, for a two-player game with ordered signals at the last stage, $r=|\mathfrak{r}|$, the feature vector has a dimension of 3. In this case, $f_{i}^{(r),0}(\vartheta)$ denotes the number of signals in $\vartheta$ where player $i$ is ranked lower than the opponent, $f_{i}^{(r),1}(\vartheta)$ denotes the number of signals where player $i$ is ranked equally with the opponent, and $f_{i}^{(r),2}(\vartheta)$ denotes the number of signals where player $i$ is ranked higher than the opponent.
@@
The potential outcome feature of player $i$ for the signal infoset $\vartheta$ at stage $r<|\mathfrak{r}|$ represents the histogram distribution of possible abstracted signal infosets that it can transition to in the next stage. Due to the potentially large size of $\mathcal{C}^{(r+1)}$, a histogram distribution with a length of $|\mathcal{C}^{r+1}|$ can incur significant computational overhead. Fortunately, in games like Texas Hold’em, we can simplify this data structure. $D=\{0,\ldots,|D|-1\}$ represents a deck of cards. In each game stage, the dealer sequentially deals a certain number of community and private cards from this deck without replacement. We define $\mathcal{U}=(\mathcal{U}^{1},\ldots,\mathcal{U}^{|\mathfrak{r}|})$ as a vector indicating the number of community cards dealt in each stage of a single game, and $\mathcal{K}=(\mathcal{K}^{1},\ldots,\mathcal{K}^{|\mathfrak{r}|})$ as a vector indicating the number of private cards dealt to each player in each stage of a single game. Additionally, we use $\upsilon=(\upsilon^{1},\ldots,\upsilon^{|\mathfrak{r}|})$ to denote the vector of community cards dealt in a single game, where $|\upsilon^{r}|=\mathcal{U}^{r}$. The private cards received by player $i$ in a single game are represented by $\kappa_{i}=(\kappa_{i}^{1},\ldots,\kappa_{i}^{|\mathfrak{r}|})$, where $|\kappa_{i}^{r}|=\mathcal{K}^{r}$. In the $r$-th stage, a signal $\theta$ can be described by the vector $(\upsilon^{1},\ldots,\upsilon^{r},\kappa_{1}^{1},\ldots,\kappa_{1}^{r},\ldots,\kappa_{N}^{1},\ldots,\kappa_{N}^{r})$, and the signal infoset $\vartheta_{i}(\theta)$ for player $i$ to which signal $\theta$ belongs can be described by the vector $(\upsilon^{1},\ldots,\upsilon^{r},\kappa_{i}^{1},\ldots,\kappa_{i}^{r})$. Firstly, Texas Hold’em-style games have perfect-recall, which implys that in a given stage, all signals within a signal infoset share the same predecessor signal infoset in the earlier stages. For example, during the Turn stage in a game of Texas Hold’em, the preceding signal infoset for the signal infoset $\left[\spadesuit A,\heartsuit A\middle|\clubsuit 3,\clubsuit 5,\diamondsuit Q,\diamondsuit K\right]$ can only be $\left[\spadesuit A,\heartsuit A\middle|\clubsuit 3,\clubsuit 5,\diamondsuit Q\right]$ during the Flop stage, and $\left[\spadesuit A,\heartsuit A\right]$ during the Preflop stage. Secondly, in Texas Hold’em-style games, the transition of signal infosets between different stages is a classical probability model. In stage $r$, a signal infoset transition to $n^{(r+1)}$ signal infosets in stage $r+1$, all with equal probability, where $n^{(r+1)}$ is calculated as
@@
For example, in Texas Hold’em, the signal infosets from the Preflop, Flop, and Turn stages can transition to signal infosets in the Flop, Turn, and River stages, respectively, with $\binom{50}{3}$, $\binom{47}{1}$, and $\binom{46}{1}$ possible cases. We can simplify these histograms using sparse representation. Specifically, for a signal infoset $\vartheta$ in stage $r$ that can transition to signal infosets $\vartheta_{1},\ldots,\vartheta_{n^{(r+1)}}$ in stage $r+1$, the potential outcome feature for $\vartheta$ is calculated as
$f_{i}^{(r)}(\vartheta)=\textsc{@NonlastStageFeature}(r,\vartheta,\mathcal{D}_{i}^{(r+1)}).$
Where the NonlastStageFearure operator is defined as shown in Algorithm 2.
@@
2:$r\in\{1,\dots,|\mathfrak{r}|-1\}$. Stages.
@@
4:$n^{(r+1)}\in\mathbb{N}$. The number of signal infosets that $\vartheta$ can transfer to at stage $r+1$.
@@
6:$\mathcal{D}_{i}^{(r+1)}:\mathbb{N}\mapsto\mathbb{N}$. Next stage potential outcome isomorphism map.
7:procedure NonlastStageFearure($r$, $\vartheta$, $\mathcal{D}_{i}^{(r+1)}$)
@@
10:         $\vartheta_{j}\leftarrow$ the $j$-th possible signal infoset in next stage of $\vartheta$.
@@
Algorithm 2 Non-last stage feature
Algorithm 1 and Algorithm 2 both involve an $\text{Index}_{i}$ operator, which is responsible for mapping the signal infoset $\vartheta\in\varTheta_{i}^{(r)}$ at stage $r$ to a unique integer ranging from 00 to $|\varTheta_{i}^{(r)}|-1$. In games like Texas Hold’em, an effective mapping method is to use colexicographic order (Bollobás, 1986). As described by Waugh (2013), grouped colexicographic encoding can be used to encode grouped combinations. The order of the signal infoset $\vartheta_{i}=(\upsilon^{1},\ldots,\upsilon^{r},\kappa_{i}^{1},\ldots,\kappa_{i}^{r})\in\varTheta_{i}^{(r)}$ can be represented as follows

### v2

@@ 3 Games with Ordered Signals
Figure 2 illustrates the game tree with ordered signal for Leduc Hold’em. According to rules of Leduc Hold’em rules, see Appendix D.1, all subtrees rooted at chance public nodes in the second phase are identical. For the sake of simplicity in visualization, we represent the Leduc game tree as a forest of two trees. However, it is worth noting that the game with ordered signals model defined in Definition 1 is a complete tree. We can concatenate a subtree from the first phase with five subtrees from the second phase to form the complete Leduc game tree. It is important to distinguish between two types of terminal public nodes. When $\tilde{z}$ is a terminal public node with fold, it implies $\sum_{i=1}^{N}\omega_{i}(\tilde{z})=1$, indicating that only one player remains in the game. On the other hand, when $\tilde{z}$ is a terminal public node with showdown, it means that more than one player is still participating in the game, and their payoffs need to be determined through a showdown. In Leduc Hold’em, Rhode Island Hold’em, HULHE, HUNL, and the Numeral211 Hold’em introduced in the experimental section, all terminal public nodes with fold are located in the final phase. Meanwhile, the signal set is solely dependent on the phase, implying that the terminal signal set $\tilde{\Theta}$ in these games corresponds exclusively to the signal set of the final phase. Unlike the definition in Gilpin and Sandholm [4], signals in Leduc Hold’em, as defined in Definition 1, refer to the complete set of cards currently dealt. For example, JQ in the first phase (the first player dealt a J, the second player dealt a Q) and JQK in the second phase (the first player dealt a J, the second player dealt a Q, and the community card is K) are signals in their respective phases. We can define other games with ordered signals where signal unveiling follows various probability distributions, not just the model of sampling without replacement, as we are familiar with. The player’s observation of the signal serves as the basis for classifying the signal into different signal infosets. Given a signal $\theta$ and a player $i\in\mathcal{N}$, we can define the signal infoset that player $i$ cannot distinguish related to $\theta$ as $\vartheta_{i}(\theta)=\{\theta^{\prime}\in\Theta\mid O_{i}(\theta)=O_{i}(\theta^{\prime})\land O_{pub}(\theta)=O_{pub}(\theta^{\prime})\}$ and $\vartheta_{c}(\theta)=\{\theta^{\prime}\in\Theta\mid O_{pub}(\theta^{\prime})=O_{pub}(\theta)\}$. Take $\theta=\text{JQK}$ as an example, the observations of $\theta$ are
$\left\{\begin{array}[]{ll}O_{1}(\theta)=\text{J**}\\ O_{2}(\theta)=\text{*Q*}\\ O_{pub}(\theta)=\text{**K}\end{array}\right..$ (4)
@@
JQK and JJK belong to the same signal infoset for player 1 while JQK and QQK belong to the same signal infoset for player 2. Additionally, the concept of a signal infoset space $\varTheta=(\varTheta_{1},\dots,\varTheta_{N},\varTheta_{c})$ is introduced, where $\varTheta_{i}=\{\vartheta_{i}(\theta)|\theta\in\Theta\}$ denotes the collection of signal infosets pertinent to player $i$. A game with ordered signals is a special type of imperfect-information game. We can transform a game with ordered signals into an imperfect-information game, a process illustrated in Appendix E.
@@
Compared to the definition of games with ordered signals introduced by Gilpin and Sandholm [4], Definition 1 is more generalized and concise. Specifically:
@@
    The definition of phases is more flexible. In the original definition, phases were predefined non-negative integers. In Definition 1, phases are defined based on the number of chance public nodes along a path. This allows games with ordered signals to be treated as a single tree structure, eliminating the need to split them into a forest based on chance public nodes.
@@ 4 Abstraction for Games with Ordered Signals
The model of games with ordered signals divides the description of nodes in imperfect-information games into two dimensions: the signals and the public nodes. Correspondingly, it categorizes the description of infosets in imperfect-information games into signal infosets and public nodes. The infosets in an imperfect-information game that can be traced back to a common public node will have the same action space and action history sequence. As long as the signal infosets are suitable for merging together, their corresponding infosets should also be suitable for merging into the same abstracted infoset. The game with ordered signals simplifies the process of infosets abstraction in imperfect-information games, eliminating the need for pre-filtering infosets. This also implies that infosets at the same phase will have consistent abstraction results based on the signal infosets abstraction. In other words, if $\vartheta_{i}^{1}$ and $\vartheta_{i}^{2}$ are two $r$-phase signal infosets for player $i$ and can be grouped into the same abstracted signal infoset during a abstraction process, then for any public node $\tilde{h}_{i}$ corresponding to player $i$ in the $r$-phase, the infoset related to $\tilde{h}_{i}$ and $\vartheta_{i}^{1}$ will always be classified into the same abstracted infoset as the one related to $\tilde{h}_{i}$ and $\vartheta_{i}^{2}$. This simplifies the workload of infoset abstraction, even though it sacrifices some flexibility in infosets abstraction, this constraint is reasonable. We introduce the following definition to provide a clearer description of the abstraction process in games with ordered signals:
Definition 2.
@@
Intuitively, one might expect that more refined abstractions would result in superior strategies. However, the discovery of abstraction pathology refutes this notion. [^2] It’s worth noting that while Waugh’s experiments focus on imperfect-information games, Leduc Hold’em itself is a form of game with ordered signals. Moreover, the information set abstraction and action abstraction performed in their experiments correspond to the signal and action abstraction of games with ordered signals , respectively. Therefore, the concept of abstraction pathology also applies to the games with ordered signals. It is crucial to stress that the issue of abstraction pathology arises from abstracting the opponent’s decision space, essentially modeling the opponent’s behavior based on a fundamental assumption: that the opponent exclusively acts within the defined abstracted decision space. However, this assumption may not always hold true. A more reasonable perspective called one-player-abstraction-perspective is that players make decisions within the abstracted decision space as a simplification to adapt to the opponent’s decisions in the complete decision space. The study conducted by Waugh et al. [14] demonstrated that when the opponent’s decision space is not abstracted, increasing the precision of one’s own abstraction can lead to reduced exploitability of the abstracted Nash equilibrium solution in the original game. [^3] It’s important to observe that although Waugh’s conclusion is derived within the context of imperfect-information games, while games with ordered signals represent a particular subclass of imperfect-information games. Consequently, the same conclusion remains valid when applied to games with ordered signals.
@@
Although the task of signal abstraction in games with ordered signals were not explicitly defined in prior research, earlier studies employed various rudimentary terms to describe similar problems and methods. This was particularly evident in tasks such as hand abstraction in the domain of Texas Hold’em game. For instance, Shi and Littman [5] referred to it as bins, Billings et al. [6] as buckets, Johanson et al. [11] mentioned state-space abstraction, and Gilpin and Sandholm [4] used the term signals to describe analogous concepts. Building upon these analogous descriptions of signal abstraction tasks, numerous approaches have been proposed to solve it. Among them, Gilpin and Sandholm [4] and Waugh [35] developed lossless isomorphism (LI), Johanson [36] introduced the expected hand strength (E[HS]) algorithm, while Gilpin et al. [9], Gilpin and Sandholm [10] introduced the potential-aware (PA) algorithm, which later led to the derivation of the potential-aware based on the earthmover’s distance (PAEMD) algorithm introduced by Ganzfried and Sandholm [12]. Our definition of the signal abstraction task has been significantly influenced by the work of Gilpin and Sandholm [4] and is also more closely aligned with the essence of the hand abstraction task in Texas Hold’em-style game.
@@ 5 Potential Outcome Isomorphism
In game theory, calculating the expected payoff at each node relies on the utility backpropagated from the terminal nodes. Consequently, mainstream signal abstraction algorithms cluster signals based on the showdown outcomes in the terminal phase, such as E[HS], PA, and PAEMD. These algorithms are referred to as outcome-based imperfect-recall algorithms. They possess the flexibility to adjust the scale of abstracted infosets. Our inquiry pertains to the resolution boundary, which refers to the quantity of signal infosets that the algorithm can recognize without exceeding, of such outcome-based imperfect-recall signal abstraction algorithms and whether they can achieve LI. The objective of this section is to develop a signal abstraction, termed as potential outcome isomorphism (POI), which aims to identify as many abstracted signal infosets as possible based on future showdown outcomes only. In the next section, we will elucidate why the POI can represent the resolution boundary of outcome-based imperfect-recall algorithms.
@@
2:$r\in\{1,\dots,|\mathfrak{r}|\}$. Phases.
@@
5:$\mathcal{D}_{i}^{(r+1)}:\mathbb{N}\mapsto\mathbb{N}$. Next phase potential outcome isomorphism map, if $r=|\mathfrak{r}|$, $\mathcal{D}_{i}^{(r+1)}=\emptyset$.
6:procedure PhasePotentialOutcomeIsomorphism($r$, $\varTheta_{i}^{(r)}$, $\mathcal{D}_{i}^{(r+1)}$)
@@
The construction of POI, as shown in Algorithm 1, is described as follows. The algorithm begins at the last phase, $r=|\mathfrak{r}|$, where $\mathfrak{r}=\{r=\gamma(\tilde{x})\mid\tilde{x}\in\tilde{X}\}$, and proceeds bottom-up, computing each phase down to $r=1$. For any player $i$ and signal information set $\vartheta$ at phase $r$, we construct a potential outcome feature $f_{i}^{(r)}(\vartheta)$. These feature vectors consist of natural numbers and have varying dimensions across phases but remain consistent within the same phase.
After removing duplicates from the potential outcome features of all signal infosets for player $i$ at phase $r$, we form a set denoted as $\mathcal{C}_{i}^{(r)}$. Upon sorting the potential outcome features in $\mathcal{C}_{i}^{(r)}$ in lexicographical order, a unique lexicographical identifier in the range of 00 to $|\mathcal{C}_{i}^{(r)}|-1$ is assigned to each potential outcome feature $f_{i}^{(r)}$. This identifier is referred to as the lexicographical order identifier ($lexid$) for $f_{i}^{(r)}$. If two signal information sets, $\vartheta_{1}$ and $\vartheta_{2}$, in $\varTheta_{i}^{(r)}$ share the same potential outcome feature $f_{i}^{(r)}$, they are categorized as the same abstracted signal infoset.
For player $i$ at the last phase, $r=|\mathfrak{r}|$, the potential outcome feature for a signal infoset $\vartheta\in\varTheta_{i}^{(r)}$ is defined as
@@
$f_{i}^{(r),l}(\vartheta)=\left\{\begin{array}[]{ll}\sum\limits_{\theta\in\vartheta}\mathds{1}\left\{\exists j\neq i,\succeq(\theta,i,j)=\text{False and }\succeq(\theta,j,i)=\text{True}\right\}&\text{if }l=0,\\ \sum\limits_{\theta\in\vartheta}\mathds{1}\left\{\sum\limits_{j\neq i}\mathds{1}\left\{\succeq(\theta,j,i)=\text{True}\right\}=N-l\text{ and }\forall j\neq i,\succeq(\theta,i,j)=\text{True}\right\}&\text{if }l\neq 0.\end{array}\right.$ (8)
@@
  • $f_{i}^{(r),l}(\vartheta)$ represents the number of signals in $\vartheta$ where player $i$ is ranked no less than all other players and is ranked exactly $l-1$ times higher than the other players.
@@
Specifically, for a two-player game with ordered signals at the last phase, $r=|\mathfrak{r}|$, the feature vector has a dimension of 3. In this case, $f_{i}^{(r),0}(\vartheta)$ denotes the number of signals in $\vartheta$ where player $i$ is ranked lower than the opponent, $f_{i}^{(r),1}(\vartheta)$ denotes the number of signals where player $i$ is ranked equally with the opponent, and $f_{i}^{(r),2}(\vartheta)$ denotes the number of signals where player $i$ is ranked higher than the opponent.
@@
The potential outcome feature of player $i$ for the signal infoset $\vartheta$ at phase $r<|\mathfrak{r}|$ represents the histogram distribution of possible abstracted signal infosets that it can transition to in the next phase. Due to the potentially large size of $\mathcal{C}^{(r+1)}$, a histogram distribution with a length of $|\mathcal{C}^{r+1}|$ can incur significant computational overhead. Fortunately, in games like Texas Hold’em, we can simplify this data structure. $D=\{0,\ldots,|D|-1\}$ represents a deck of cards. In each game phase, the dealer sequentially deals a certain number of community and private cards from this deck without replacement. We define $\mathcal{U}=(\mathcal{U}^{1},\ldots,\mathcal{U}^{|\mathfrak{r}|})$ as a vector indicating the number of community cards dealt in each phase of a single game, and $\mathcal{K}=(\mathcal{K}^{1},\ldots,\mathcal{K}^{|\mathfrak{r}|})$ as a vector indicating the number of private cards dealt to each player in each phase of a single game. Additionally, we use $\upsilon=(\upsilon^{1},\ldots,\upsilon^{|\mathfrak{r}|})$ to denote the vector of community cards dealt in a single game, where $|\upsilon^{r}|=\mathcal{U}^{r}$. The private cards received by player $i$ in a single game are represented by $\kappa_{i}=(\kappa_{i}^{1},\ldots,\kappa_{i}^{|\mathfrak{r}|})$, where $|\kappa_{i}^{r}|=\mathcal{K}^{r}$. In the $r$-th phase, a signal $\theta$ can be described by the vector $(\upsilon^{1},\ldots,\upsilon^{r},\kappa_{1}^{1},\ldots,\kappa_{1}^{r},\ldots,\kappa_{N}^{1},\ldots,\kappa_{N}^{r})$, and the signal infoset $\vartheta_{i}(\theta)$ for player $i$ to which signal $\theta$ belongs can be described by the vector $(\upsilon^{1},\ldots,\upsilon^{r},\kappa_{i}^{1},\ldots,\kappa_{i}^{r})$. Firstly, Texas Hold’em-style games have perfect-recall, which implys that in a given phase, all signals within a signal infoset share the same predecessor signal infoset in the earlier phases. For example, during the Turn phase in a game of Texas Hold’em, the preceding signal infoset for the signal infoset $\left[\spadesuit A,\heartsuit A\middle|\clubsuit 3,\clubsuit 5,\diamondsuit Q,\diamondsuit K\right]$ can only be $\left[\spadesuit A,\heartsuit A\middle|\clubsuit 3,\clubsuit 5,\diamondsuit Q\right]$ during the Flop phase, and $\left[\spadesuit A,\heartsuit A\right]$ during the Preflop phase. Secondly, in Texas Hold’em-style games, the transition of signal infosets between different phases is a classical probability model. In phase $r$, a signal infoset transition to $n^{(r+1)}$ signal infosets in phase $r+1$, all with equal probability, where $n^{(r+1)}$ is calculated as
@@
For example, in Texas Hold’em, the signal infosets from the Preflop, Flop, and Turn phases can transition to signal infosets in the Flop, Turn, and River phases, respectively, with $\binom{50}{3}$, $\binom{47}{1}$, and $\binom{46}{1}$ possible cases. We can simplify these histograms using sparse representation. Specifically, for a signal infoset $\vartheta$ in phase $r$ that can transition to signal infosets $\vartheta_{1},\ldots,\vartheta_{n^{(r+1)}}$ in phase $r+1$, the potential outcome feature for $\vartheta$ is calculated as
$f_{i}^{(r)}(\vartheta)=\textsc{@NonlastPhaseFeature}(r,\vartheta,\mathcal{D}_{i}^{(r+1)}).$
Where the NonlastPhaseFearure operator is defined as shown in Algorithm 2.
@@
2:$r\in\{1,\dots,|\mathfrak{r}|-1\}$. Phases.
@@
4:$n^{(r+1)}\in\mathbb{N}$. The number of signal infosets that $\vartheta$ can transfer to at phase $r+1$.
@@
6:$\mathcal{D}_{i}^{(r+1)}:\mathbb{N}\mapsto\mathbb{N}$. Next phase potential outcome isomorphism map.
7:procedure NonlastPhaseFearure($r$, $\vartheta$, $\mathcal{D}_{i}^{(r+1)}$)
@@
10:         $\vartheta_{j}\leftarrow$ the $j$-th possible signal infoset in next phase of $\vartheta$.
@@
Algorithm 2 Non-last phase feature
Algorithm 1 and Algorithm 2 both involve an $\text{Index}_{i}$ operator, which is responsible for mapping the signal infoset $\vartheta\in\varTheta_{i}^{(r)}$ at phase $r$ to a unique integer ranging from 00 to $|\varTheta_{i}^{(r)}|-1$. In games like Texas Hold’em, an effective mapping method is to use colexicographic order [37]. As described by Waugh [35], grouped colexicographic encoding can be used to encode grouped combinations. The order of the signal infoset $\vartheta_{i}=(\upsilon^{1},\ldots,\upsilon^{r},\kappa_{i}^{1},\ldots,\kappa_{i}^{r})\in\varTheta_{i}^{(r)}$ can be represented as follows